OpenStudy (anonymous):

a) what number is 1/4 of the way from 16 to 256? b) what common fraction is one-third of the way from 1/4 to 1/3?

hero (hero):

256 - 16 = 240 240/4 = 60 60 + 16 = 76 1/3 - 1/4 = 4/12 - 3/12 = 1/12 1/12 divided by 3 = 1/36 1/4 + 1/36 = 9/36 + 1/36 = 10/36 = 5/9 So the answer is a = 76 b = 5/9

OpenStudy (anonymous):

oh thanks a million

hero (hero):

These are not that easy. You can easily make a mistake with these

OpenStudy (anonymous):

true on that